# MN-TSG: Continuous Time Series Generation with Irregular Observations

### v1

@@ 4.1 Experimental Settings
Here we briefly outline the experimental settings, with details shown in the Appendix E.1.
Datasets. The datasets we used include: (1) four popular public time series generation datasets, namely Sines, Stocks, Energy, and MuJoCo. (2) Medical ECG-related time series classification datasets with a variety of different temporal patterns from the UCR archive, including ECG200, ECG5K, ECGFD, and TLECG. (3) Synthetic datasets, including integrated datasets of different signal waves and polynomial coefficient fitting datasets.
Baselines. We compare our method with several representative advanced time series generation methods, including KO-VAE (Naiman et al., 2024), GTGAN (Jeon et al., 2022), ProFITi Yalavarthi et al. (2025) and HeTVAE (Shukla and Marlin, ) tailored for irregular time series generation, as well as TimeGAN (Yoon et al., 2019), TimeVAE (Desai et al., 2021), and diffusion model (Huang et al., 2025) used for regular TSG.
Implementation details. We conduct experiments using the publicly available baseline code and train it to convergence with the recommended settings. In our method, the number of experts is fixed at four, and the standard diffusion model is adopted as the generation backbone. For the four public datasets and two synthetic datasets, we construct samples for training and evaluation using three different lengths: 12, 24, and 36, and finally report the average metrics. For the medical ECG classification datasets, each sample has already been well processed, and we use the default parameters. We randomly discard 30%, 50%, and 70% of the temporal observations along the time dimension to construct irregularly sampled datasets (Naiman et al., 2024).
4.2 Performance of Irregular Time Series Generation (TSG)
Quantitative evaluation. In this experiment, we assess our model’s performance on irregular time series generation. We integrate regular time series generation methods with standard NCDE, using it to impute missing values before generation. Results in Tables 1 and 2 show that our method consistently outperforms baselines in both popular TSG and medical datasets. While baselines handle irregular data via NCDE, their performance is limited due to optimization challenges and difficulties in capturing dynamic patterns. In contrast, our MOE-NCDE uses a mixture of experts and a decoupled training strategy with a pre-trained channel-wise autoencoder to focus on learning MOE-dynamics. This makes our method better capture dynamic temporal patterns and more effective for irregular time series data.
@@ 4.2 Performance of Irregular Time Series Generation (TSG)
Qualitative evaluation. We also qualitatively assess the similarity between real and generated sequences using two visualizations: (i) t-SNE projection into two dimensions, and (ii) kernel density estimation of their probability density functions (PDFs). As shown in Figure 3, in the t-SNE plots (top), our approach shows high overlap between real and synthetic samples, while our PDFs (bottom) display more similar trends than baselines KO-VAE and GT-GAN.
Privacy protection analysis. We adopt the Membership Inference Risk (MIR) metric (Tian et al., 2024) to assess whether synthetic data generated by various models reveals information about the real data, as shown in Appendix Table 10 due to limited space. While our method achieves superior fidelity in terms of DS, KL, and MDD, the MIR results also show that our model exhibits the lowest privacy leakage risk among all methods. This indicates that our jointly generated dynamics and data not only improve quality but also show lower privacy leakage risk.
4.3 Performance of Continuous Time Series Generation (TSG)
Refined continuous generation on the downstream task. In this experiment, we evaluate the benefits of continuous TSG with irregular observations. Specifically, we begin by defining a multivariate prediction task. Baselines use regular generation with length data to predict the last time step. Then, we apply NCDE for continuous generation, in which we generate new data points in the middle of every two time steps and insert them into the original window. As a result, the length of the historical window is approximately doubled, called refined sample. Finally, we further use the refined sample to predict the same last time step to compare the performance with baselines. We set and report average MSE and MAE. Since KOVAE and GTGAN are not designed for continuous generation, we use their trained NCDE to interpolate generated samples for refinement.
Table 3 shows that the performance of KOVAE-NCDE and GTGAN-NCDE is even worse than that of their non-continuous generation counterparts KOVAE and GTGAN, indicating that the method of using standard NCDE for continuous generation is not effective. In contrast, Ours- is better than its non-continuous version, improving the forecasting performance. This is attributed to our carefully designed MOE-NCDE and the joint training process, which not only generates new samples but also produces MOE-weights tailored for each sample, thereby enabling better continuous generation tasks using the MOE-NCDE. The visualization analysis of why continuous time series is effective can be found in Appendix F.2.
@@ 4.3 Performance of Continuous Time Series Generation (TSG)
Refined continuous generation on solving for the analytical solutions of polynomial functions. We further assess continuous generation via analytical solutions using a cubic function . For each coefficient , 2500 samples are drawn from the normal distributions with different means and standard deviations. With linearly spaced over 24 values in , we compute to form time series (visualized in Appendix figure 28). Then, the model is trained on , aiming for coefficients recalculated from generated to match the original distributions.
Figure 4 shows that analytical solutions from our generated series closely match the true distribution. Both irregular to regular generation and irregular to continuous generation preserve distributional similarity, with refined continuous generation (Fig. 4(d)) closer to ground truth than non-continuous. This confirms the effectiveness of our approach, whereas baselines relying on standard NCDE deviate significantly (Appendix Figures 29, 30 and 31).
4.4 Ablation Study
Based on the experiments of imputing irregular time series data, we demonstrate the effectiveness of each component of MOE-NCDE from both quantitative and qualitative perspectives.
Quantitative evaluation. Table 4 shows that removing MOE dynamics and the decoupled design, the model’s accuracy decreased to varying degrees. When both of them are removed, the performance further deteriorates, indicating that the proposed MOE dynamics and decoupled design are crucial components. The former helps NCDE learn dynamic patterns, while the latter decouples the training of MOE-NCDE by using a pre-trained channel-wise autoencoder to replace the and the reconstruction function in NCDE, thereby better focusing on the learning of the MOE-dynamics.
@@ 4.4 Ablation Study
| w/o Decoupled | 0.370 | 0.191 | 0.878 | 0.397 | 0.174 | 0.850 | 0.361 | 0.154 | 0.865 | |
| w/o MOE-Decoup. | 0.412 | 0.173 | 0.965 | 0.363 | 0.175 | 0.968 | 0.361 | 0.157 | 0.984 | |
@@
| w/o Decoupled | 0.466 | 0.094 | 0.035 | 0.502 | 0.081 | 0.028 | 0.836 | 0.083 | 0.034 | |
| w/o MOE-Decoup. | 0.877 | 0.109 | 1.203 | 0.583 | 0.096 | 1.211 | 0.667 | 0.074 | 1.323 | |
Qualitative evaluation. We construct synthetic signals (e.g., sine, piecewise, sawtooth) and divide them into training and test sets. We train the MOE-NCDE and observe its generation effect on the test set under component ablations (Fig. 5). Only using both MOE dynamics and decoupled design does the model produce data resembling the real signals, while removing either causes high errors, confirming their effectiveness in our decoupled MOE Neural CDE.
Comparisons between Mixture-of-experts and single-expert parameter stacking. Additional experiments confirm the benefit of cooperation among multiple experts. Table 5, compares NCDE performance between (i) four experts (each with one MLP and two linear layers, eight layers in total) and (ii) a single expert with four MLPs stacked (same eight layers). The mixture-of-experts approach significantly outperforms the single expert with stacked layers, demonstrating that having mixture-of-experts enables better learning of diverse temporal dynamics.
@@
Visualizing specific temporal dynamics learned by MoE-NCDE. Figure 6 demonstrates the strategy of decomposing complex temporal dynamics for separate learning and then integrating them, which substantially enhances NCDE’s capability to model diverse temporal behaviors.
4.5 Hyperparameter Analysis and Time Complexity Analysis
@@
Hyperparameter Analysis. Figure 7 shows that MOE-NCDE performance is not highly sensitive to the number of experts: using more than one can both improve performance, with four experts yielding satisfactory performance. Appendix Figure 26 shows the analysis on the DS metric, which also draws a similar conclusion. However, the default four experts may not be the optimal parameter, indicating that there is still room for further improvement in our metrics.
@@
Time Complexity Analysis. Naturally, MOE-NCDE has higher time complexity than standard NCDE (e.g., 1.2s vs. 0.81s for double-length refined generation on 3,674 stock samples of length 12) due to mixture-of-experts dynamic functions. However, since we learn the joint distribution of the experts’ weights and time series samples, the model not only generates samples but also assigns appropriate weights to them simultaneously. This avoids the need to re-extract features and perform forward propagation to compute weights, thereby reducing computational overhead.
5 Conclusion and Limitations
This work investigates the task of continuous time series generation. On the one hand, to enhance NCDE’s ability to impute data at arbitrary time intervals, we propose a mixture-of-experts dynamic function and a decoupled design, called Decoupled MOE Neural CDE. This helps NCDE better focus on learning the dynamic functions, thereby enabling it to subsequently better handle irregular data and continuous generation tasks. On the other hand, we offer new insights into the continuous generation task by proposing to learn the joint distribution of dynamic functions and time series samples. This allows the model to generate new samples while assigning them appropriate dynamic function weights for better continuous time series generation. Extensive experiments on ten datasets show that our method outperforms advanced baselines in both quantitative and qualitative evaluations for different generation tasks.
Limitations. We have an assumption that the underlying dynamics are continuous and differentiable, which lies in the requirements of Neural CDE. This smoothness assumption enables the use of differential equation solvers for trajectory modeling but restricts their applicability in certain real-world scenarios where dynamics often involve abrupt changes, discontinuities, or non-differentiable behaviors. In future works, we may examine potentials to resolve or relax this assumption.
Appendix A Appendix
A.1 Related Work
Time Series Generation for regular Data.
@@ A.1 Related Work
Time series generation has gained increasing attention as a way to produce synthetic data for training, privacy preservation, and simulation. Early work explored adversarial learning approaches such as TimeGAN [27], which leverage adversarial objectives to jointly capture temporal dynamics and sample realism. Variational methods have also been developed, including TimeVAE [4] and Vector Quantized TimeVAE [10]. More recently, diffusion-based models [14] have emerged, synthesizing high-fidelity sequences by iteratively denoising random noise. These approaches illustrate the rapid methodological progress in generative modeling for time series.
Time Series Modeling with Irregular Data.
@@
Irregularly sampled data becomes a critical challenge in time series modeling. An important line of work introduces continuous-time dynamics to naturally handle uneven sampling, such as Neural ODEs [2] and their extensions, which enrich the dynamics with additional inputs like the slope of imputed observations [9], also called Neural CDE. Other approaches directly encode timestamp information into sequence models, e.g., using specialized irregularity-aware architectures [28]. However, it cannot model the temporal patterns at arbitrary time points and cannot be directly applied to continuous generation tasks. Collectively, these methods underscore the importance of explicitly modeling temporal irregularity as a prerequisite for generating reliable and useful sequences.
Based on Neural CDE, several methods for handling irregular-to-regular data generation have been proposed, such as those based on VAE [16] and GAN [8]. However, these methods haven’t explored how to improve NCDE, and haven’t investigated how to perform and evaluate continuous time series generation tasks. In contrast, this paper conducts an in-depth study and discussion on these aspects.
Appendix B Channel-wise autoencoder and its connection to NCDE
Explanation of pre-training part: We pre-train an MLP-based encoder ()–decoder () pair on the original irregular data by performing channel-wise encoding and reconstruction. All non-channel dimensions are flattened, and the model is trained via MSE loss to minimize the reconstruction error. Since the input and output dimensions of and match the number of time-series variables (channels), they can be directly used to map hidden states within NCDE. The full channel-wise autoencoder procedure is now given as Algorithm 4 in the revised manuscript.
Explanation of z(0): The latent initial state represents the encoded hidden representation of the time series at . The latent trajectory follows the controlled differential equation (Eq. 2). To obtain , we take the observed values at from and map them into the latent space using the pre-trained encoder .
Explanation of reconstruction function: After solving the NCDE, we obtain latent trajectories, . To recover the projected variables, we apply the decoder , yielding , which serves as the reconstruction function.
Appendix C Preliminary
Neural ordinary differential equations
@@ Appendix C Preliminary
Neural ODEs [2] provide a continuous-time modeling framework particularly suitable for irregularly-sampled time series. The latent dynamics are governed by the following differential equation:
@@
where is the latent state at time , and is a neural network parameterized by represent the dynamics of the differential equations. The latent state can be initialized as . is an encoder that maps the observed time series to the latent space. A decoder can be used to map the latent trajectory back to the data space to reconstruct or generate the continuous signal.
Appendix D More discussions about contributions
The core novelty of our method lies not in “simple combination" but in synergistic innovations tailored to the unaddressed challenge of irregular-to-continuous time series generation, which requires rethinking how these components interact to solve domain-specific pain points.
First, our adaptation of MoE to NCDE is novel in both purpose and mechanism. Unlike conventional MoE applications in NLP or computer vision, we explicitly design the MoE dynamics to capture heterogeneous temporal behaviors (e.g., diverse ECG rhythms across patients with different diseases) that standard NCDEs struggle to model. This is enabled by a task-specific router that maps irregular time-series characteristics—such as missing-value patterns and local volatility—into expert weights, enabling dynamic specialization across diverse temporal dynamics. Existing MoE or NCDE research has not explored this direction for time series generation. Furthermore, MoE allows us to decouple the modeling of temporal dynamics into expert-weight vectors that can be learned independently and generated through diffusion, ensuring that every newly generated sample receives a tailored dynamic function.
Second, we introduce a decoupled training strategy that addresses a known limitation of NCDE optimization. Prior NCDE-based models optimize both the initial latent mapping and the dynamic function jointly, often leading to optimization conflicts (e.g., reconstruction dominating dynamics learning). Our pre-trained channel-wise autoencoder provides a clean and stable initialization for and reconstruction, allowing NCDE to focus specifically on learning the underlying temporal dynamics. This design is essential to making MoE-NCDE not only feasible but also significantly more effective.
Finally, we extend diffusion models beyond standard TSG generation. Existing diffusion-based TSG works only model the distribution of observed values. We propose a joint-generation mechanism that simultaneously generates observations and MoE weights, ensuring that each synthesized time series is coupled with its own specialized dynamic function for downstream continuous refinement. This closes the methodological gap between “regular sequence generation” and “continuous generation,” a capability not previously addressed in the literature.
In summary, we propose the irregular-to-regular-to-continuous generation framework, MN-TSG, in which we improve NCDE so it can be seamlessly coupled with any state-of-the-art regular time-series generator to learn their joint distribution. Afterwards, MN-TSG injects the newly sampled NCDE-MOE weights into the pre-trained NCDE, enabling fine-grained and continuous generation for each newly generated TS data.
Appendix E Additional experimental settings
E.1 Datasets

### v2

@@ 4.1 Experimental settings
Datasets. The datasets include: (1) four public TSG datasets, Sines, Stocks, Energy, and MuJoCo; (2) ECG-related medical TS datasets from the UCR archive with diverse temporal patterns, including ECG200, ECG5K, ECGFD, and TLECG, with 2, 5,2 and 2 classes respectively; and (3) synthetic datasets, consisting of integrated signal waves and polynomial coefficient fitting data. For irregular modeling, we randomly discard 30%, 50%, and 70% of temporal observations to construct irregularly sampled datasets (Naiman et al., 2024). Details are provided in Appendix D.1.
Baselines. We compare our method with several representative advanced TSG methods, including irregular TSG methods: KO-VAE (Naiman et al., 2024), GTGAN (Jeon et al., 2022), ProFITi Yalavarthi et al. (2025) and HeTVAE (Shukla and Marlin, ), as well as regular TSG methods: TimeGAN (Yoon et al., 2019), TimeVAE (Desai et al., 2021), and diffusion model (Huang et al., 2025). For regular TSG methods, we utilize standard NCDE to impute missing values before generation.
Implementation details. We conduct experiments using publicly available baseline code and train all models to convergence with recommended settings. In our method, the number of experts is fixed at four. More parameter settings, e.g., the diffusion model, are provided in the Appendix D.2.
For the four public datasets and two synthetic datasets, we construct training and evaluation samples with three sequence lengths (12, 24, and 36) and report averaged metrics. For the medical ECG classification datasets, samples are officially preprocessed, and default parameters are used.
Evaluation metrics. Following previous works (Naiman et al., 2024; Huang et al., 2025), we apply the following metrics for evaluating generation performance: (1) Discriminative score (DS), (2) Marginal distribution difference (MDD), (3) Kullback-Leibler divergence (KL), (4) Membership Inference Risk (MIR) metric (Tian et al., 2024) for privacy protection evaluation.
4.2 Performance of irregular-to-regular TSG
Quantitative evaluation. We first evaluate our model on the irregular TSG task. As shown in Tables 1 and 2, our method consistently outperforms baselines on both popular TSG and medical datasets across various missing rate settings. Although baselines handle irregular data via NCDE, their performance is limited by weak dynamic modeling and coupled optimization design. In contrast, our MoE-NCDE leverages a mixture-of-experts function and a dynamics-focused optimization strategy with a pre-trained channel-wise autoencoder, enabling more effective learning of temporal dynamics for irregular TS.
@@ 4.2 Performance of irregular-to-regular TSG
Qualitative evaluation. We also qualitatively evaluate the similarity between real and generated sequences using two visualizations: (i) t-SNE projections and (ii) kernel density estimation of probability density functions (PDFs). As shown in Figure 3, the t-SNE plots (top) exhibit strong overlap between real and synthetic samples, while the PDFs (bottom) show closer alignment with real data compared to KO-VAE and GT-GAN.
Privacy protection analysis. This experiment evaluates whether synthetic data generated by different models leaks information from real data, with results reported in Appendix Table 9. Our method not only achieves superior fidelity in DS, KL, and MDD but also exhibits the lowest privacy leakage risk according to MIR. This is due to the MoE design, which promotes diverse generated samples rather than mere replication of the original data.
4.3 Performance of irregular-to-continuous TSG
Refined continuous TSG on the downstream task. In this experiment, we evaluate continuous TSG with irregular observations using a multivariate prediction task, which is conducted on the newly generated samples of each method. Baselines predict the last time step from length data. For Diff-MN, we apply MoE-NCDE for newly generated samples and MoE weights from the diffusion model. For KOVAE and GTGAN, we apply the trained standard NCDE to their newly generated samples for Continuous TSG. Continuous TSG generates new points between each time step, creating a refined sample that roughly doubles the length of the historical window. Following previous downstream evaluations Naiman et al. (2024); Jeon et al. (2022), the refined sample is then used to train the Gated Recurrent Unit (GRU) to predict the same last step. We set and report average MSE and MAE.
Table 3 shows that KOVAE-NCDE and GTGAN-NCDE perform worse than their non-continuous counterparts, indicating that standard NCDE for continuous generation is ineffective. In contrast, Ours- outperforms its non-continuous version, improving forecasting performance. This is due to our diffusion parameterized MoE-NCDE, which generates MoE weights capturing the temporal dynamics of the newly generated samples, enabling better continuous generation. Visualizations of continuous TS are in Appendix Figures 10 to 12.
@@ 4.3 Performance of irregular-to-continuous TSG
Refined continuous TSG on solving for the analytical solutions of polynomial functions. We further assess continuous TSG using a cubic function . For each coefficient, 2500 samples are drawn from normal distributions with varying means and standard deviations. With linearly spaced over 24 values in , forms time series (visualized in Appendix figure 16). The model is trained on , aiming for coefficients derived from generated to match the original distributions.
Figure 4 shows that analytical solutions from our generated closely match the true distribution. Both irregular-to-regular and irregular-to-continuous TSG preserve distributional similarity, with refined continuous generation (Fig. 4(d)) closer to ground truth than non-continuous. This confirms the effectiveness of our method. In contrast, baselines using standard NCDE deviate significantly (Appendix Figures 17, 19, 20).
@@
| w/o Decoupled Design | 0.370 | 0.191 | 0.878 | 0.397 | 0.174 | 0.850 | 0.361 | 0.154 | 0.865 | |
| w/o MoE + w/o Decoupled Design | 0.412 | 0.173 | 0.965 | 0.363 | 0.175 | 0.968 | 0.361 | 0.157 | 0.984 | |
@@
| w/o Decoupled Design | 0.466 | 0.094 | 0.035 | 0.502 | 0.081 | 0.028 | 0.836 | 0.083 | 0.034 | |
| w/o MoE + w/o Decoupled Design | 0.877 | 0.109 | 1.203 | 0.583 | 0.096 | 1.211 | 0.667 | 0.074 | 1.323 | |
@@
4.4 Ablation study
Quantitative evaluation. Table 4 shows that removing MoE functions or the decoupled design significantly decreases model accuracy, and removing them together further deteriorates performance, indicating that these components are crucial. The MoE functions help NCDE learn temporal patterns, while the dynamics-focused decoupled optimization design (called Decoupled Design) uses a pre-trained channel-wise autoencoder to replace the network SIN and RN, allowing better focus on learning MoE dynamics.
Qualitative evaluation. We construct synthetic signals (e.g., sine, piecewise, sawtooth) and split them into training and test sets. We train MoE-NCDE and evaluate generation on the test set under component ablations (Fig. 5). Only with both MoE dynamics and the dynamics-focused decoupled optimization design does the model produce signals resembling the real data; removing either causes high errors, confirming their effectiveness.
Exploring why MoE is effective. Table 5 compares NCDE performance between (i) four experts (each with one MLP and two linear layers, eight layers total) and (ii) a single expert with four stacked MLPs (same eight layers). The MoE approach significantly outperforms the single expert, as each expert specializes in learning distinct temporal dynamics, making the overall dynamics easier to learn while enhancing both expressiveness and generalization. Figure 6 further illustrates how MoE decomposes complex temporal patterns across experts and then integrates them, substantially enhancing NCDE’s ability to model diverse temporal behaviors. Moreover, Appendix Table 13 reports the average expert weights across datasets. The weights are non-sparse, and different datasets favor different experts, indicating that the Mixture of Experts structure is effectively utilized rather than collapsing to a single expert.
4.5 Hyperparameter and time complexity analysis
@@
Hyperparameter analysis. Figure 7 shows that MoE-NCDE performance is not highly sensitive to the number of experts: using more than one improves performance, with four experts yielding satisfactory results.
@@
Time complexity analysis. MoE-NCDE has higher time complexity than standard NCDE (e.g., 1.2s vs. 0.81s for double-length refined generation on 3,674 stock samples of length 12) due to mixture-of-experts dynamics. However, by parameterizing the MoE weights via the diffusion model, we can directly generate sample-specific MoE weights for new data, avoiding NCDE retraining, additional feature extraction, and forward propagation, thereby reducing computational overhead.
5 Conclusion
This paper presents the first study of continuous TSG. Built upon NCDEs, Diff-MN introduces a decoupled MoE-NCDE that enhances standard NCDEs with mixture-of-experts dynamic functions and dynamics-focused optimization design, enabling more effective modeling of complex temporal dynamics and irregular TS for continuous generation. Moreover, we parameterize MoE weights via a diffusion model to generate sample-specific weights that reflect their temporal dynamics. Combined with a pre-trained NCDE, this enables adaptive continuous TSG. Diff-MN achieves state-of-the-art performance on continuous generation across multiple public and synthetic datasets.
Appendix A Supplementary related work
TSG for regular data.
@@ Appendix A Supplementary related work
TSG has gained increasing attention as a way to produce synthetic data for training, privacy preservation, and simulation. Early work explored adversarial learning approaches such as TimeGAN [30], which leverage adversarial objectives to jointly capture temporal dynamics and sample realism. Variational methods have also been developed, including TimeVAE [6] and Vector Quantized TimeVAE [12]. More recently, diffusion-based models [17] have emerged, synthesizing high-fidelity sequences by iteratively denoising random noise. These approaches illustrate the rapid methodological progress in generative modeling for TS. Recently, an increasing number of generation methods based on diffusion models have been proposed [4, 5, 9, 28, 14], such as task-oriented generation, cross-domain generation, causal generation, and text-controlled generation.
TS modeling with irregular data.
@@
Irregularly sampled data becomes a critical challenge in TS modeling. An important line of work introduces continuous-time dynamics to naturally handle uneven sampling, such as Neural ODEs [2] and their extensions, which enrich the dynamics with additional inputs like the slope of imputed observations [11], also called Neural CDE. Other approaches directly encode timestamp information into sequence models, e.g., using specialized irregularity-aware architectures [31]. However, it cannot model the temporal patterns at arbitrary time points and cannot be directly applied to continuous generation tasks. Collectively, these methods underscore the importance of explicitly modeling temporal irregularity as a prerequisite for generating reliable and useful sequences.
Appendix B Discussion about why using neural controlled differential equation (NCDE)
B.1 Comparisons between NODE and NCDE
B.1.1 Theoretical perspective
@@
Neural ordinary differential equation (NODE) [2] provides a continuous-time modeling framework for irregularly-sampled time series. The latent dynamics are governed by the following differential equation:
@@
where is the latent state at time , and is a neural network parameterized by represent the dynamics of the differential equations.
However, NODE evolves the latent state solely from based on learned dynamics, without direct conditioning on the observed values. Consequently, the latent trajectory may drift away from the true data distribution.
We adopt NCDE because Diff-MN requires both imputing irregular observations and performing continuous generation. Values generated between existing observation points need to be consistent with the temporal dynamics of the surrounding context; otherwise, the generated values may deviate from the true distribution of the underlying dynamical system.
B.1.2 Experimental perspective
GT-GAN incorporates ODE-RNN, and ablation studies of GT-GAN consistently show that it outperforms the two ODE-based methods. Additionally, KoVAE (a TSG method using NCDE) is reported to perform even better than GT-GAN, and our Diff-MN outperforms both GT-GAN and KoVAE. All of this indicates that NCDE-based TSG under observed-data constraints outperforms ODE-based methods.
B.2 Proving interpolation control in NCDE does not cause over-smoothing
B.2.1 Theoretical perspective
As discussed in the main text, NCDE can be defined as
where is the control path (obtained via interpolation and smoothing), and is a nonlinear function (neural network).
The neural network can model highly nonlinear dynamics, amplifying or bending small variations in the smooth control path and thereby inducing complex, rapidly changing behavior in the latent trajectory.
Hence, even if is smooth, the behavior of the integral result depends on the properties of . If is highly nonlinear, the derivative can vary rapidly with , leading to non-smooth trajectories.
Specifically, when is smooth and differentiable, the NCDE admits the following differential form:
Although the interpolated control path is smooth by construction, the evolution of is governed by the nonlinear function . As a result, even when is smooth, the product may vary rapidly due to the nonlinearity of , leading to non-smooth latent trajectories.
In our model, the is further parameterized as a Mixture-of-Experts:
where and are routing weights produced by the router , and denotes the -th expert.
Substituting the MoE formulation into the NCDE dynamics yields
Since both the expert functions and the routing weights depend nonlinearly on the latent state, different experts may dominate the dynamics at different time instances. This allows the latent trajectory to exhibit rapid local variations and highly diverse temporal behaviors.
Therefore, even with a smooth interpolated control path , the NCDE equipped with MoE dynamics can generate expressive and potentially non-smooth latent trajectories.
B.2.2 Experimental perspective
To validate this, we added visualizations of sequences generated solely by NCDE, as shown Figure 13, 14 and 15. These results show: (i) For intrinsically smooth data (e.g., sine in Figure 16), NCDE produces smooth trajectories and (ii) For intrinsically smooth data (e.g., sine in Figure 16), NCDE produces smooth trajectories.
Overall, while the control path is smooth, the nonlinear MoE–NCDE dynamics allow the generated sequences to exhibit fluctuating temporal behaviors, depending entirely on the underlying temporal dynamics.
Appendix C Supplementary algorithm pseudocodes
This presents the algorithm pseudocodes omitted from the Method section of the main text due to space constraints. As all algorithms have already been properly referenced and discussed, we simply include them here for convenient reference.
C.1 Decoupled optimization for MoE-dynamics-focused training (Algorithm 2)
C.2 Training channel-wise autoencoder (Algorithm 1)
C.3 Joint diffusion training for to parameterize MoE weights (Algorithm 3)
C.4 Continuous time series generation (Algorithm 4)
Appendix D Supplementary experimental settings
D.1 Datasets